OpenStudy (anonymous):

what is the absolute value of 2x+1 greater than or equal to -1 ?

OpenStudy (anonymous):

first of all this is not possible. The absolute value can not equal a negative number but: same or opposite Same opposite 2x+1>=-1 -(2x+1)>=-1 2x>=0 2x+1<=1 x>=0 2x<=0 x>=0 or x<=0 which is all real numbers but there is no solution to this problem!

OpenStudy (anonymous):

thank you! is it no solution or all reals?

OpenStudy (anonymous):

actually I mis stated it, since it is greater than or equal to, all real numbers are greater than -1. If it was only =, we would get 0 but that would not be a valid solution. so all real numbers